I need to know how to calculate the surface area of a circular cone that has a radius of 3 and height of 7. I have tried it many times and I get stuck calculating the square root portion can you help me?

Answer by stanbon(57307)

I need to know how to calculate the surface area of a circular cone that has a radius of 3 and height of 7.
--------------
surface area = pi*r^2 + prL where L is the slant height of the cone.
---------------------------------
L = sqrt(3^2+7^2) = sqrt(58) = 7.62
----
SA = (3.14)*3^2 + (3.14)*3*7.62 = 100 sq. units
==================================================
Cheers,
Stan H.